# Resolving Build Conflicts via Example-Based and Rule-Based Program Transformations

### v1

@@ Abstract.
Merge conflicts often arise when developers integrate changes from different software branches. The conflicts can result from overlapping edits in programs (i.e., textual conflicts), or cause build and test errors (i.e., build and test conflicts). They degrade software quality and hinder programmer productivity. While several tools detect build conflicts, few offer meaningful support for resolvinntional cases like those caused by method removal. To overcome limitations of existing tools, we introduce BuCoR (BUild COnflict Resolver), a new conflict resolver. BuCoR first detects conflicts by comparing three versions related to a merging scenario: base $b$, left $l$, and right $r$. To resolve conflicts, it employs two complementary strategies: example-based transformation (BuCoR-E) and rule-based transformation (BuCoR-R). BuCoR-R applies predefined rules to handle common, well-understood conflicts. BuCoR-E mines branch versions ($l$ and $r$) for exemplar edits applied to fix related build errors. From these examples, it infers and generalizes program transformation patterns to resolve more complex conflicts.
We evaluated BuCoR on 88 real-world build conflicts spanning 21 distinct conflict types. BuCoR generated at least one solution for 65 cases, and correctly resolved 43 conflicts. We observed that this hybrid approach—combining context-aware, example-based learning with structured, rule-based resolution—can effectively help resolve conflicts. Our research sheds light on future directions of more intelligent and automated merge tools.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software evolution
@@ 1. Introduction
As illustrated in Figure 1, a typical merging scenario involves five program versions: two branch versions $l$ and $r$ whose edits need to be merged, base version $b$—the common origin of both branches, the automatically merged version $A_{m}$ produced by git-merge when it naïvely integrates branch edits textually, and the manually merged version $m$ that developers create based on $A_{m}$. Among the various possible conflicts between $l$ and $r$, build conflicts refer to the incompatible edits whose naïve integration triggers build errors in the resulting merged version. As shown in Figure 3, because $l$ adds a call to $m()$ while $r$ renames that method, the co-application of both edits can cause a compilation error of unresolved method reference. Namely, the newly added method call $m()$ is not associated with any defined method.
Figure 1. A typical merging scenario can involve up to five program versions
Although tools were created to detect build conflicts (Towqir et al., 2023; Wuensche et al., 2020a; Shen et al., 2019), there is rare automatic support to resolve such conflicts. In particular, Gmerge (Zhang et al., 2022) relies on Clang compiler message to locate conflict-related changes in branches, and applies k-shot learning with GPT-3 to suggest symbol renaming for conflict resolution. MrgBldBrkFixer (Sung et al., 2020) starts with a naïvely merged version $A_{m}$, looks for build errors due to failed resolution of symbols, identifies symbol renaming changes in branches-to-merge, extracts related patches in either branch, and similarly applies those patches to fix build conflicts. These tools suffer from three limitations:
@@
Figure 3. BuCoR leverages Bucond (Towqir et al., 2023) to detect build conflicts, and employs two complimentary strategies to resolve conflicts
  1. (1)
    Both tools adopt compiler errors to locate conflicts; if compilers fail to locate errors, neither tool works.
@@
To overcome those limitations, we introduce BuCoR (BUild COnflict Resolver), a novel resolver of build conflicts. As shown in Figure 3, given the three program versions ($l$, $b$, $r$) related to a merging scenario, BuCoR first applies Bucond (Towqir et al., 2023) to reveal conflicting edits between branches. As Bucond detects conflicts via static analysis instead of automatic build, BuCoR eliminates dependencies on compilers. BuCoR also adopts git-merge to generate a naïvely merged version $A_{m}$, extending Bucond to (1) create a graphical representation for $A_{m}$ and (2) relate that graph with those of the three input program versions. Based on the generated conflict report and four-way interconnected graph, BuCoR opportunistically applies two complementary strategies: BuCoR-E and BuCoR-R.
@@
Rule-based transformation (BuCoR-R): Prior studies reveal patterns frequently applied to resolve certain kinds of build conflicts (Shen et al., 2022; Da Silva et al., 2022). Following those studies, we defined and implemented 16 resolution rules/patterns in BuCoR-R. Given a conflict, BuCoR-R searches its rule set for an applicable rule; when a rule is found, BuCoR-R customizes that rule for edit application.
To evaluate the tool effectiveness, we applied BuCoR to 88 real-world build conflicts in 30 open-source projects. BuCoR correctly resolved 43 conflicts. BuCoR-E and BuCoR-R separately generated resolutions for 28 and 51 conflicts; 21 and 29 of those resolutions separately match developers’ resolutions recorded in version history. All these numbers demonstrate BuCoR’s effectiveness in handling conflicts. To sum up, we made the following contributions:
  • We created BuCoR, a new static analysis-based resolver of build conflicts, to novelly combine example-based transformation with rule-based transformation. Unlike prior work, BuCoR does not adopt compilers to detect conflicts.
@@
Changes in $\boldsymbol{l}$ (responsible def updates): In TypeSerializerConfig.java, names of the file, Java class, and constructor are all updated to SerializerConfig. Changes in $\boldsymbol{r}$ (responsible use introduction): In a newly added file XmlClientConfigBuilder.java, a method is defined to access TypeSerializerConfig + private void handleSerializers(Node node, SerializationConfig serializationConfig) { + + if (”type-serializer”.equals(name)) { + TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); + typeSerializerConfig.setClassName(value); + final String typeClassName = getAttribute(child, ”type-class”); + typeSerializerConfig.setTypeClassName(typeClassName); + serializationConfig.addTypeSerializer(typeSerializerConfig); + } } Edit example in $\boldsymbol{l}$ to adapt usage of TypeSerializerConfig: private void handleSerializers(Node node, …) { - if(”type-serializer”.equals(name)){ - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); - typeSerializerConfig.setClassName(value); + if(”serializer”.equals(name)){ + SerializerConfig serializerConfig = new SerializerConfig(); + serializerConfig.setClassName(value); final String typeClassName=retrieveAttribute(child, ”type-class”); - typeSerializerConfig.setTypeClassName(typeClassName); - serializationConfig.addTypeSerializer(typeSerializerConfig); + serializerConfig.setTypeClassName(typeClassName); + serializationConfig.addSerializerConfig(serializerConfig);
Tool-generated edits that are applicable to $A_{m}$, to resolve conflicts
(c) A naïve resolution producible by existing tools and BuCoR-R (d) A more comprehensive resolution produced by BuCoR-E
private void handleSerializers(Node node, …) { if (”type-serializer”.equals(name)) { -   TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); +   SerializerConfig typeSerializerConfig = new SerializerConfig(); typeSerializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); typeSerializerConfig.setTypeClassName(typeClassName); serializationConfig.addTypeSerializer(typeSerializerConfig); } } private void handleSerializers(Node node, …) { - if (”type-serializer”.equals(name2)) { -  TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); -  typeSerializerConfig.setClassName(value); + if (”serializer”.equals(name2)) { +  SerializerConfig serializerConfig = new SerializerConfig(); +  serializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); -  typeSerializerConfig.setTypeClassName(typeClassName); -  serializationConfig.addTypeSerializer(typeSerializerConfig); +  serializerConfig.setTypeClassName(typeClassName); +  serializationConfig.addSerializerConfig(serializerConfig); } }
@@ 2. A Motivating Example
To facilitate discussion, this section introduces a running example we crafted based on a real-world project hazelcast (haz, 2022). As shown in Table 1(a), a merging scenario has $l$ rename a class and $r$ insert code to use the original class. The naïve integration of these edits can trigger a build error as the newly introduced uses refer to a nonexistent class def; thus, a build conflict occurs.
To resolve such a conflict, existing tools simply update class references by replacing TypeSerializerConfig with SerializerConfig (Table 1(c)). However, such a replacement is insufficient, as the context still has variable typeSerializerConfig and literal “type-serializer” match the original class usage. Consequently, after existing tools update class references, developers need to manually replace defs and uses of related variables/literals, to ensure consistent updates and prevent semantic conflicts. Furthermore, when such class uses are introduced at multiple places, developers have to go over all places to manually apply those edits again and again, which process is tedious and error-prone.
To overcome the limitations of existing tools and better help developers, we introduce BuCoR—a hybrid approach to combine example-based resolution with rule-based resolution. The example-based resolution BuCoR-E is derived from our insight, on the association between build errors in branches and build conflicts in the merged version. Basically, many build errors are caused by mismatches between def and use of the same program entities; many build conflicts are caused by mismatches between modified def and newly introduced use of entities (Shen et al., 2022; Da Silva et al., 2022). If on either branch, developers resolved def-use mismatches by applying specialized edits; then they are likely to reapply the same or similar edits to resolve conflicts that show the same kind of mismatches in software merge.
@@
For the responsible use-introduction of each conflict, BuCoR-E tentatively establishes context matching between $P$ and the edit location. If the matching succeeds, BuCoR-E customizes $P$ by generating edit operations with respect to the matched context. BuCoR then applies those operations to transform code. As shown in Table 1(d), the context-to-handle is different from the original edit context (see Table 1(b)): it uses a different variable (name2 vs. name), and invokes a different method (getAttribute(...) vs. retrieveAttribute(...)). Albeit the differences, BuCoR still resolves the conflict by mimicking developers’ coding practices.
@@ 3. Approach
[…]
For implementation, Bucond uses JavaParser to parse source code, and adopts JGraphT (jgr, 2025) to construct and analyze PEGs.
@@ 3.3. Example-based Transformation (BuCoR-E)
Given a conflict, BuCoR-E identifies the responsible def-change, mines the branch contributing that change, and looks for any exemplar edit which adapts entity uses and surrounding code for that change. In our motivating example, because $l$ renames class TypeSerializerConfig and its constructor, it is the contributor branch of def-change. By traversing entity-level edits of $l$$\Delta(G_{l},G_{b})$, BuCoR-E locates changes of the corresponding class/constructor references. Namely, if an entity $\boldsymbol{e}$ uses TypeSerializerConfig in $b$ but uses SerializerConfig in $l$, then the entity $e$ contains an edit example. Certainly, if multiple entities adapt their usage to the same def-change, BuCoR-E considers all these entities to have relevant examples.
@@
[…]
@@ 3.3.3. Phase III: Pattern Application
/* Find a match for the last usage of $e$/
@@ 0.14        Search for a best match in a forward manner
 /* Compare parent nodes of $s_{p}$ and $s_{m}$/
@@ 0.21  Check whether $p_{m}$ matches $p_{p}$
As illustrated by Algorithm 1, given a pattern $P$, a method $M$ whose entity use causes a conflict, and the changed entity $e$, BuCoR-E tentatively matches nodes in a bottom-up manner with the best effort. Specifically, it first locates a statement $s_{p}$ in $P$ that contains the last usage of $e$, to search among $M$ for a statement $s_{m}$ best matching $s_{p}$ (see 1.1–1.2). To compare statements, we check their AST node types and values to calculate a matching score as below:
@@
  • Value Match: Two nodes, $x$ and $y$, have a value match if their strings have a 3-gram similarity score greater than 0.618. Here, the string of a simple statement (e.g, MethodInvocation) includes all content of that statement; the string of a complex statement (e.g., conditional or loop structure) only includes the structure header (e.g., if-condition or for-loop header). A successful value match increments the overall matching score by the calculated similarity value. This threshold-based approach allows similar nodes to match, even though they are not identical. Such a flexibility also increases the applicability of inferred patterns.
@@
Once a best match of $s_{p}$ is found (with score $>$ 1.618), BuCoR-E treats the pair $(s_{p},s_{m})$ as anchor points to guide node matching for their siblings (see 1.5–1.14). Specifically for siblings preceding $s_{p}$, it enumerates nodes in reverse order: it first compares $s_{p}$’s nearest sibling against all siblings preceding $s_{m}$ for a best match; it then proceeds to match $s_{p}$’s more distant siblings against any remaining unmatched nodes, moving backward from the most recently matched one. Similarly, for siblings following $s_{p}$, BuCoR-E matches nodes in forward order.
After searching best matches for all sibling nodes of $s_{p}$, the process continues to tentatively match the parent node of $s_{p}$ with that of $s_{m}$. If this match succeeds, BuCoR-E repeats 1.3–1.17 in Algorithm 1, using the newly matched blocks as anchors for further match. The procedure continues until no more parent node in $P$ can be found or the parent match fails. In the end, the algorithm outputs best matches of nodes, and their matching scores.
If Phase II infers multiple patterns from edit examples, these patterns may be all applicable to the method-to-edit $M$. To find the best one, we defined two ways to score the overall context matching between each pattern and $M$, based on the output of Algorithm 1:
@@
We believe that the greater the contextual similarity between a candidate pattern $P_{i}$ and $M$, the easier it is to apply that pattern and the more likely the resulting version is correct. Thus, among all applicable patterns, we select the one with the highest $\Sigma_{m}$. If multiple candidates have a tie, we choose the one with the highest $C_{m}$, ensuring a closer semantic match between that pattern and $M$.
In our motivating example, the method-to-edit (Table 1(c)) shares a highly similar context with the inferred pattern (Listing 1): $\Sigma_{m}=9.94$ and $C_{m}=4$. All five edited statements in the pattern sequentially find best matches within the method: four statements match exactly, and the if-condition has a highly similar counterpart.

### v2

# Combining Example-Based and Rule-Based Program Transformations to Resolve Build Conflicts

@@ Abstract.
Merge conflicts often arise when developers integrate changes from different software branches. The conflicts can result from overlapping edits in programs (i.e., textual conflicts), or cause build and test errors (i.e., build and test conflicts). They degrade software quality and hinder programmer productivity. While several tools detect build conflicts, few offer meaningful support for resolving them. To overcome limitations of existing tools, we introduce BuCoR (BUild COnflict Resolver), a new conflict resolver. BuCoR first detects conflicts by comparing three versions related to a merging scenario: base $b$, left $l$, and right $r$. To resolve conflicts, it employs two complementary strategies: example-based transformation (BuCoR-E) and rule-based transformation (BuCoR-R). BuCoR-R applies predefined rules to resolve conflicts in frequently suggested or conventional ways. BuCoR-E mines branch versions ($l$ and $r$) for exemplar edits applied to fix related build errors. From these examples, it infers and generalizes program transformation patterns to resolve conflicts in project-specific or unconventional ways.
We evaluated BuCoR on 88 real-world build conflicts spanning 21 distinct conflict types. BuCoR generated at least one solution for 65 cases, and correctly resolved 34 conflicts. We observed that this hybrid approach—combining context-aware, example-based learning with structured, rule-based resolution—can effectively help resolve conflicts. Our research sheds light on future directions of more intelligent and automated merge tools.
@@
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793347isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Software maintenance toolsccs: Software and its engineering Maintaining softwareccs: Software and its engineering Software evolution
@@ 1. Introduction
As illustrated in Figure 1, a typical merging scenario involves five program versions: two branch versions $l$ and $r$ whose edits need to be merged, base version bb—the common origin of both branches, the automatically merged version $A_{m}$ produced by git-merge when it naïvely integrates branch edits textually, and the manually merged version $m$ that developers create based on $A_{m}$. Among the various possible conflicts between $l$ and $r$, build conflicts refer to the incompatible edits whose naïve integration triggers build errors in the resulting merged version. As shown in Figure 3, because $l$ adds a call to $m()$ while $r$ renames that method, the co-application of both edits can cause a compilation error of unresolved method reference. Namely, the newly added method call $m()$ is not associated with any defined method.
Figure 1. A merging scenario involves five program versions
@@
Figure 3. BuCoR leverages Bucond (Towqir et al., 2023) to detect build conflicts, and employs two complementary strategies to resolve conflicts
Although tools were created to detect build conflicts (Towqir et al., 2023; Wuensche et al., 2020), there is rare automatic support to resolve such conflicts. In particular, Gmerge (Zhang et al., 2022) relies on Clang compiler messages to locate conflict-related changes in branches, and applies k-shot learning with GPT-3 to suggest symbol renaming for conflict resolution. MrgBldBrkFixer (Sung et al., 2020) starts with a naïvely merged version $A_{m}$, looks for build errors due to failed resolution of symbols, identifies symbol renaming changes in branches-to-merge, extracts related patches in either branch, and similarly applies those patches to fix build conflicts. These tools suffer from two limitations:
@@
To overcome these limitations, we introduce BuCoR (BUild COnflict Resolver), a novel resolver of build conflicts. As shown in Figure 3, given the three program versions ($l$, $b$, $r$) related to a merging scenario, BuCoR first applies Bucond (Towqir et al., 2023), a static analysis-based tool, to reveal conflicting edits between branches. BuCoR also adopts git-merge to generate a naïvely merged version $A_{m}$, extending Bucond to (1) create a graphical representation for $A_{m}$ and (2) relate that graph with those of the three input program versions. Based on the generated conflict report and four-way interconnected graph, BuCoR opportunistically applies two complementary strategies: BuCoR-E and BuCoR-R.
@@
Rule-based transformation (BuCoR-R): Prior studies reveal patterns frequently applied to resolve certain kinds of build conflicts (Wuensche et al., 2020; Sung et al., 2020). Following those studies, we defined and implemented 16 resolution rules/patterns in BuCoR-R. Given a conflict, BuCoR-R searches its rule set for an applicable rule; when a rule is found, BuCoR-R customizes that rule for edit application.
To evaluate the tool effectiveness, we applied BuCoR to 88 real-world build conflicts in 30 open-source projects. BuCoR correctly resolved 34 conflicts. BuCoR-E and BuCoR-R separately generated resolutions for 28 and 51 conflicts; 21 and 20 of those resolutions separately match developers’ resolutions recorded in version history. All these numbers demonstrate BuCoR’s effectiveness in handling conflicts. To sum up, we made the following contributions:
  • We created BuCoR, a new static analysis-based resolver of build conflicts, to novelly combine example-based transformation with rule-based transformation.
@@
At https://github.com/shadabtowqir/BuCoR, we open-sourced our program and data.
@@
Changes in $\boldsymbol{l}$ (responsible def updates): In TypeSerializerConfig.java, names of the file, Java class, and constructor are all updated to SerializerConfig. Changes in $\boldsymbol{r}$ (responsible use introduction): In a newly added file XmlClientConfigBuilder.java, a method is defined to access TypeSerializerConfig + private void handleSerializers(Node node, SerializationConfig serializationConfig) { + + if (”type-serializer”.equals(name)) { + TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); + typeSerializerConfig.setClassName(value); + final String typeClassName = getAttribute(child, ”type-class”); + typeSerializerConfig.setTypeClassName(typeClassName); + serializationConfig.addTypeSerializer(typeSerializerConfig); + } }      Edit example in $\boldsymbol{l}$ to adapt usage of TypeSerializerConfig: private void handleSerializers(Node node, …) { - if(”type-serializer”.equals(name)){ - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); - typeSerializerConfig.setClassName(value); + if(”serializer”.equals(name)){ + SerializerConfig serializerConfig = new SerializerConfig(); + serializerConfig.setClassName(value); final String typeClassName=retrieveAttribute(child, ”type-class”); - typeSerializerConfig.setTypeClassName(typeClassName); - serializationConfig.addTypeSerializer(typeSerializerConfig); + serializerConfig.setTypeClassName(typeClassName); + serializationConfig.addSerializerConfig(serializerConfig);    Tool-generated edits that are applicable to $A_{m}$, to resolve conflicts    (c) A naïve resolution producible by existing tools and BuCoR-R (d) A more comprehensive resolution produced by BuCoR-E private void handleSerializers(Node node, …) { if (”type-serializer”.equals(name)) { - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); + SerializerConfig typeSerializerConfig = new SerializerConfig(); typeSerializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); typeSerializerConfig.setTypeClassName(typeClassName); serializationConfig.addTypeSerializer(typeSerializerConfig); } } - if (”type-serializer”.equals(name2)) { - TypeSerializerConfig typeSerializerConfig = new TypeSerializerConfig(); - typeSerializerConfig.setClassName(value); + if (”serializer”.equals(name2)) { + SerializerConfig serializerConfig = new SerializerConfig(); + serializerConfig.setClassName(value); final String typeClassName=getAttribute(child, ”type-class”); - typeSerializerConfig.setTypeClassName(typeClassName); - serializationConfig.addTypeSerializer(typeSerializerConfig); + serializerConfig.setTypeClassName(typeClassName); + serializationConfig.addSerializerConfig(serializerConfig); } }
@@ 2. A Motivating Example
To facilitate discussion, we introduce a running example we crafted based on a real-world project hazelcast (haz, 2022). As shown in Table 1(a), a merging scenario has $l$ rename a class TypeSerializerConfig and $r$ insert code to use the original class. The naïve integration of these edits can trigger a build error as the newly introduced uses refer to a nonexistent class def; thus, a build conflict occurs.
To resolve such conflicts, existing tools update class references by replacing TypeSerializerConfig with SerializerConfig (Table 1(c)). However, such a replacement is insufficient, as the context still has variable typeSerializerConfig and literal “type-serializer” match the original class. Consequently, after existing tools update class references, developers need to manually replace defs and uses of related variables/literals, to ensure consistent updates and prevent semantic conflicts. When such class uses are introduced at multiple places, developers have to go over all places to manually apply those edits again and again, which process is tedious and error-prone.
To overcome the limitations of existing tools and better help developers, we introduce BuCoR—a hybrid approach to combine example-based resolution with rule-based resolution. The example-based resolution BuCoR-E is derived from our insight, on the association between build errors in branches and build conflicts in the merged version. Basically, many build errors are caused by mismatches between def and use of the same program entities. There is commonality between (1) fixes to build errors on branches, and (2) resolutions to build conflicts in the merged version  (Shen et al., 2022; Da Silva et al., 2022). If on either branch, developers resolved def-use mismatches by applying specialized edits; then they are likely to reapply similar edits to resolve conflicts that show the same kind of mismatches in software merge.
@@
For the responsible use-introduction of each conflict, BuCoR-E establishes context matching between $P$ and the edit location. If a match is found, BuCoR-E customizes $P$ by creating edit operations with respect to the matched context. BuCoR then applies those operations to transform code. As shown in Table 1(d), the context-to-handle is different from the original edit context (see Table 1(b)): it uses a different variable (name2 vs. name), and calls a different method (getAttribute(...) vs. retrieveAttribute(...)). However, BuCoR could resolve the conflict by mimicking developers’ edits.
@@ 3. Approach
Bucond statically analyzes three program versions related to each merging scenario—$b$, $l$, rr—to identify build conflicts. It models each version as a program entity graph (PEG), which captures defined program entities (e.g., classes and methods) and their relations (e.g., type reference or method calls). By comparing PEGs, Bucond extracts entity-level edits in $l$ and $r$. It then matches these edits with 57 predefined patterns of conflicting edits. For instance, one pattern checks when a method is renamed in one branch, whether the other branch adds any call to the original method. Bucond reports a conflict if any match is found. Unlike compiler-based tools, Bucond does not try to generate or build the naïvely merged version $A_{m}$. It can detect conflicts even if $A_{m}$ is uncompilable, and pinpoint the specific edits responsible—which compilers cannot do.
For implementation, Bucond uses JavaParser (jp, 2025) to parse source code, and adopts JGraphT (jgr, 2025) to construct and analyze PEGs.
@@ 3.3. Example-based Transformation (BuCoR-E)
[…]
@@
  • update($t$, $v_{n}$): To replace the old value of node $t$ with the new value $v_{n}$.
  • add($t$, $t_{p}$, $i$, $l$, $v$): To add a new node $t$ to the AST, as the $i^{th}$ child of node $t_{p}$. Here, $l$ and $v$ separately specify $t$’s entity type (e.g., method invocation) and its value (e.g., the statement string).
  • delete($t$): To remove a node $t$ from the AST.
  • move($t$, $t_{p}$, $i$): To move a node $t$ and its subtree to the $i^{th}$ child of node $t_{p}$.
@@ 3.3.3. Phase III: Pattern Application
/* Find a match for the last usage of $e$ */
@@ 0.14        Search for a best match in a forward manner
 /* Compare parent nodes of $s_{p}$ and $s_{m}$ */
@@ 0.21  Check whether $p_{m}$ matches $p_{p}$
As shown in Algorithm 1, given a pattern $P$, a method $M$ whose entity use causes a conflict, and the changed entity $e$, BuCoR-E performs best-effort bottom-up node matching. It first locates in $P$ a statement $s_{p}$ that contains the last usage of $e$, and searches $M$ for a statement $s_{m}$ best matching $s_{p}$ (lines 1.1–1.2).
To compare statements, we check their AST node types and values to calculate a matching score as below:
@@
  • Value Match: Two nodes, $x$ and $y$, have a value match if their strings have a 3-gram similarity score greater than 0.618; the similarity score is added to the total matching score. The string of a simple statement (e.g., MethodInvocation) includes all content of that statement; the string of a complex statement (e.g., loop) only includes the header. This threshold-based approach allows flexible matching across similar code, and increases the applicability of inferred patterns.
@@
Once a best match of $s_{p}$ is found (with score >> 1.618), BuCoR-E treats the pair $(s_{p},s_{m})$ as anchor points to guide sibling match (lines 1.5–1.14). Preceding siblings are matched in reverse order, starting from the nearest sibling, while following siblings are matched in forward order. The 1.618 threshold reflects the expectation of both a type match and a value similarity of at least 0.618. Here, 0.618 is the golden ratio used in prior work (Shen et al., 2019) to yield robust matching.
After matching all siblings, BuCoR-E attempts to match the parent nodes of $s_{p}$ and $s_{m}$. If this match succeeds, BuCoR-E repeats 1.3–1.17 in Algorithm 1, using the newly matched blocks as anchors for further match. The procedure continues until no further parent matches are possible. In the end, the algorithm outputs best matches of nodes, and their matching scores.
If multiple patterns inferred from Phase II apply to method $M$, BuCoR-E ranks them with two metrics derived from Algorithm 1:
@@
We believe that the greater the contextual similarity between a candidate pattern $P_{i}$ and $M$, the easier it is to apply that pattern and the more likely the resulting version is correct. Thus, among all applicable patterns, we select the one with the highest $\Sigma_{m}$. Ties are broken using $C_{m}$ to favor closer semantic matches.
In our motivating example, the method-to-edit (Table 1(c)) shares a highly similar context with the inferred pattern (Listing 1): $\Sigma_{m}=9.94$ and $C_{m}=4$. All five edited statements in the pattern sequentially find best matches within the method: with four exact matches and a highly similar if-condition.